OpenStudy (anonymous):

PLEASE HELP!!!!!!!! Evaluate the following limit. lim as h->0 of 1/h integral from 5 to 5+h of sqrt(7+t^3)dt. It said Hint: Use L'Hopital's rule. I will write this as an equation below in the next box as well.

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0}\frac{ 1 }{ h}\int\limits_{5}^{5+h}\sqrt{7+t^3}dt\]

OpenStudy (anonymous):

this is exactly F'(5)=sq(7+5*3)